OpenStudy (anonymous):

What is the value of 4(x+2)=14-2(3-4x)

OpenStudy (anonymous):

hi

OpenStudy (photon336):

On the left hand side we can see here that this is the distributive property \[4(x+2) = 4*2+4*2 = 4x+8 \] now let's work on the right hand side we make sure we distribute the negative \[14-2(3-4x) = 14-(2*3)-(2*)(-4x) = 14-6+8x\] \[4x+8 = 8+8x \] \[4x = 0 , x = 0\] we get x = 0 so we check our answer \[4(0+2) = 14-2(3-0), 8 = 8 \]

OpenStudy (photon336):

@skyepilot17

OpenStudy (photon336):

Does this make sense?

OpenStudy (anonymous):

@Photon336 yes, Thank you!

OpenStudy (photon336):

The key here is the distributive property \[a(b+c) = a*b + a*c = ab+ac \]